# Spectrum Sharing in Satellite-Terrestrial Integrated Networks: Frameworks, Approaches, and Opportunities

### v1

@@ Abstract
To accommodate the increasing communication needs in non-terrestrial networks (NTNs), wireless users in remote areas may require access to more spectrum than is currently allocated. Terrestrial networks (TNs), such as cellular networks, are deployed in specific areas, but many underused licensed spectrum bands remain in remote areas. Therefore, bringing NTNs to a shared spectrum with TNs can improve network capacity under reasonable interference management. However, in satellite-terrestrial integrated networks (STINs), the comprehensive coverage of a satellite and the unbalanced communication resources of STINs make it challenging to effectively manage mutual interference between NTN and TN. This article presents the fundamentals and prospects of spectrum sharing (SS) in STINs by introducing four SS frameworks, their potential application scenarios, and technical challenges. Furthermore, advanced SS approaches related to interference management in STINs and performance metrics of SS in STINs are introduced. Moreover, a preliminary performance evaluation showcases the potential for sharing the spectrum between NTN and TN. Finally, future research opportunities for SS in STINs are discussed.
@@ I Introduction
Due to the boosting need for broader and more reliable wireless network coverage and data transmission, non-terrestrial networks (NTNs) have emerged as a promising solution. In urban areas with heavy traffic or remote areas where terrestrial networks (TNs) face geographical deployment limitations, NTNs can provide seamless access and ubiquitous connectivity. Spectrum integration in satellite-terrestrial integrated networks (STINs) is a necessary step to improve the utilization efficiency of scarce spectrum resources. However, the coexistence of NTNs and TNs may be challenging, especially with regard to spectrum utilization and interference management. When NTN nodes such as low-earth orbit (LEO) satellites are incorporated into existing TNs via certain frequency bands, interference patterns can become more complicated due to different spatial propagation characteristics of signals. For example, the use of the Ka band or the Ku band for satellites may cause severe co-channel interference in existing communication systems. Moreover, within a fixed spectrum allocation, it is possible that while certain portions of the spectrum in NTNs are fully occupied, other segments within TNs may remain unutilized. The imbalance will lead to under-utilization of available spectrum resources and a significant hindrance to substantial improvements in system performance.
@@
In recent years, the rapid growth of NTNs has paved the way toward realizing STINs. STINs, a promising candidate for future 6th-generation (6G) wireless communications, incorporate LEO nodes into existing TN to take advantage of NTN and meet diverse communication requirements [1, 2]. By using STINs, the coverage, flexibility, and spectrum efficiency of the wireless network can be significantly promoted, as terrestrial users can access various TN and NTN nodes from vast geographical areas. Inevitably, the spectrum resource will become scarcer in STIN as more NTN nodes are introduced. This encourages us to manage the spectrum when TN and NTN coexist in the STIN [3]. Meanwhile, the licensed spectrum to the static TN nodes cannot always be utilized efficiently, as spatial spectrum holes cannot be eliminated within the geographic area of interest. Hence, mobile NTN users can admit access to users within the spectrum holes and promote spectrum efficiency [4]. In existing work, one of the promising technologies for TN and NTN coexistence is spectrum sharing (SS), which allows the NTN to share the spectrum pre-allocated to the TN. SS allows the NTN to work in the same frequency band when the TN users are active, as long as its interference to the TN can be tolerated [4, 5]. In this way, the STIN can provide high-quality and seamless wireless services at a slight cost of performance loss under the practical restrictions of limited spectrum resources.
@@
With the distinct advantages above, STINs can overwhelm conventional TNs in a variety of practical fields. As the key technique in STINs, SS becomes the main bottleneck in network performance promotion and faces several challenges [2]. First, a universal SS architecture cannot always meet the diverse needs of the STIN. Therefore, the SS frameworks for all possible STIN transmission scenarios should be carefully designed, including the uplink and downlink of both TN and NTN. Second, since interference in STIN leads to system performance degradation, mitigating interference becomes the most critical issue. Generally, such mitigation can be implemented by isolating users from the interference source and carefully managing the TN and NTN spectrum [6, 7]. Meanwhile, the TN and NTN spectrum should be allocated appropriately, and the spectrum access scheme should be carefully designed [8]. In addition, since STINs are hierarchical and heterogeneous, the proper selection of performance metrics for system evaluation is also essential to the system, depending on the design objectives.
@@
To address the above concerns, we first consider the isolation of STIN users. One of the well-known approaches for this is the utilization of a protection zone for TN, where NTN users within the protection zone are not allowed to access the spectrum shared by TN and NTN [6]. However, the appropriate definition of protection zones still remains a challenging problem. Meanwhile, some studies have been carried out for spectrum allocation and management. Typical work involves spatial SS, dynamic SS, and game-theoretic SS approaches, which are potentially beneficial for STINs. In spatial SS, the objective is to find the geographic areas that are good for NTN access in STINs [9]. In dynamic SS, the objective usually becomes finding the temporal holes in the spectrum resource by treating the NTN as a secondary network [3, 10]. Both spatial SS and dynamic SS can be considered as frequency reuse schemes in STINs. For spectrum allocation and sharing, game models are competitive candidates in finding the optimal scheme by modeling the TN and NTN as game players [11, 12]. All the above topics can be further studied to meet the performance requirements of future STINs.
@@
This article focuses on the SS frameworks for various STIN scenarios and the state-of-the-art approaches to SS in STINs. Our main contributions are as follows.
@@
  • We summarize the potential performance metrics for evaluating the efficiency of SS in STINs and validate the metrics via a practical case study. Finally, future research opportunities of SS in STINs are highlighted.
The remainder of the paper is organized as follows. Section II discusses the SS frameworks for various STIN scenarios, including uplink and downlink for TN and NTN in the network architecture. In Section III, we discuss advanced approaches to SS in STINs. In Section IV, we perform practical SS case studies to validate the above frameworks for STINs. Section V summarizes the future research directions of SS in STINs. Finally, the conclusion remarks are given in Section VI, and an overview of SS in STINs is provided in Fig. 1.
@@ II Frameworks of SS in STINs
In this section, we present four frameworks of SS in STINs based on different combinations of uplink and downlink in NTN and TN, as illustrated in Fig. 2. In scenarios without SS, NTN and TN operate on their respective dedicated frequencies. However, this leads to the inefficient use of TN’s dedicated spectrum in some remote areas. By implementing SS for NTN and TN in the development of STINs, the total bandwidth is divided into two segments: 1) shared spectrum for both TN and NTN, and 2) reserved spectrum exclusively for NTN, as shown at the top of Fig. 2. This approach allows NTN and TN to share a substantial amount of spectrum resources flexibly and efficiently. Notably, a portion of the spectrum is reserved for NTN because specific existing TN equipment may not be compatible with the higher frequency bands used by NTN. Therefore, this reserved spectrum can only be utilized by NTN. The following subsections provide a detailed introduction to the four frameworks depicted in Fig. 2.
@@ II-A NTN Downlink and TN Downlink SS
When a satellite transmits signals to an NTN user and a BS transmits signals to a TN user, they can access the shared spectrum, as shown in Fig. 2(1). In this case, TN users can be regarded as primary users and NTN users as secondary users. Designing a three-dimensional (3D) protection zone for each TN user is essential. By adopting such a strategy, NTN users should ensure that there are no TN users within the radius distance of the protection zone before receiving signals. In addition, satellites cannot be present within the 3D area that includes TN users and BSs. This is because the sharing of spectrum between TN and NTN will not only enhance the communication performance of TN users by virtue of the expanded bandwidth but also result in interference. The protection zone is mainly set to protect the quality of service (QoS) of TN users. Theoretically, NTN users can also be affected by interference leakage from BSs, considering their beam directionality on the antenna surface in the horizontal plane. However, near-vertical transmissions from satellites can hardly be influenced. Moreover, satellites can mitigate mutual interference by constraining spectrum access within a specified distance range.
@@ II-B NTN Downlink and TN Uplink SS
Consider a scenario involving sharing a reverse link spectrum in which an NTN user receives signals from a satellite. As depicted in Fig. 2(2), both the satellite and the NTN user can access the shared spectrum through TN uplink communications. Satellite spectrum sensing can be performed to detect the strength of TN uplink signals to reduce mutual interference. If the uplink signal exhibits a low level of strength, the NTN downlink transmission can be operated to prevent significant mutual interference. To be more specific, due to the directionality of satellite downlink signals, the corresponding interference to TN uplink transmissions on the near-horizontal plane is relatively small. In addition, owing to the extended propagation distance and limited antenna gain associated with the satellite’s side lobes, TN uplink users located within the reception range of these side lobes exert minimal interference on the satellites’ spectrum sensing capabilities. In particular, spectrum sensing is conducted because satellites and NTN users can neither determine TN users transmitting in the current time slot nor predict aggregated interference.
@@ II-C NTN Uplink and TN Downlink SS
When an NTN user transmits signals to a satellite, they can opportunistically access the shared spectrum with TN downlink communications, as shown in Fig. 2(3). In this situation, TN users are considered primary users, while NTN users are secondary users. This framework is feasible due to a couple of factors. First, the BS is tilted downward for TN downlink communications and employs beamforming techniques to serve TN users. Because of the BS’s antenna angle and beamforming, interference leakage from the BS to the satellite is minimal. Second, during NTN uplink communications, the NTN user and the satellite can collaboratively perform spectrum sensing. This allows them to opportunistically access the spectrum while minimizing interference to and from TN users. It is important to note that, based on spectrum sensing conducted by NTN users, the aggregated interference power generated by NTN users is relatively lower than the desired signal power of the TN downlink communication. This is due to the high transmit power associated with TN BS, which also employs beamforming techniques.
@@ II-D NTN Uplink and TN Uplink SS
As shown in Fig. 2(4), allowing NTN uplink communications to share the spectrum with TN uplink communications is generally not advisable for several reasons. First, numerous TN uplink users operate within a satellite’s broad coverage area, and the distances between these users and NTN uplink users can be significant. This results in limited detection capabilities, making it difficult for either type of user to perceive one another effectively. Second, TN users typically have a limited number of antennas, leading to considerable interference leakage from TN users in the vertical direction. This interference can overwhelm the desired signals from NTN uplink communications. Third, NTN users are likely to experience a high rate of missed detections regarding the presence of TN users. Consequently, this inability to accurately sense TN users can lead to significant interference from NTN uplink transmissions to TN communications. As a result, TN users’ communication performance often falls short of expectations. Sharing the spectrum between these two co-directional uplink links is unfavorable from the perspective of interference and sensing capabilities in many situations. However, this spectrum-sharing framework may be viable with appropriate spectrum-aware strategies when the number of TN uplink users is relatively small.
@@ III-A Protection Zone-Based Approach
SS technique enables STIN users to sense and access the spectrum occupied by other users to improve the spectrum efficiency, and interference to the primary system can be tolerated when performance degradation is insignificant. However, such interference cannot be ignored when users are close to interference sources. Therefore, an auxiliary approach for SS is to define some geographical areas where users cannot access the spectrum of the primary system. Known as protection zones, these geographical regions are usually located around the primary users, which can isolate interference for the users inside them, as shown in Fig. 3(a). With protection zones and user locations in STINs, the SS technique enables intelligent monitoring of the spatial spectrum holes and shares the spectrum outside the protection zone. Since the NTN is mobile and provides flexible spectrum access for users, finding the boundary for the protection zone becomes a crucial and challenging issue. In addition, the dynamics of STIN architectures also have significant impacts on the protection zones, and we should consider them in the successful design of SS strategies.
III-B Spatial Spectrum Sensing-Based Approach
@@
Spatial spectrum sensing is an approach to accessing the spectrum that relies on sensing rather than knowing the positions of interfering nodes [13]. The term “spatial” refers to the ability of this method to capture the spatial characteristics of large-scale networks. For a secondary user seeking to access a spectrum occupied by primary users, the process begins with evaluating the interference power. As shown in Fig. 3(b), if the test statistic for the interference power exceeds a predefined energy detection threshold, it indicates that the channel is currently busy. In this case, the secondary user transmits with a low probability. Conversely, if the interference power is below the threshold, the secondary user transmits with a high probability. It is essential to ensure that the false-alarm and miss-detection probabilities are controlled following the Neyman-Pearson criterion. When spatial spectrum sensing is implemented in STINs, NTN-related devices act as secondary users to monitor the busy status of the shared channel. By selecting an appropriate energy detection threshold, we can manage the interference generated by NTN toward TN.
III-C Joint Spatial Spectrum Sensing-Based Approach
@@
Joint spatial spectrum sensing involves two different types of network nodes working together to perform spatial spectrum sensing. As shown in Fig. 3(c), if both nodes determine that the channel is available for transmission, the secondary user will likely access the channel with a high probability. Conversely, if they disagree on availability, the secondary user will have a lower probability of accessing the channel. This approach is particularly suitable for NTN uplink and TN downlink SS. The reason is that the satellite, acting as the NTN receiver, may experience significant interference from the TN, especially when there is a high density of ground base stations in specific areas. As a result, the spatial spectrum sensing performed by the satellite protects its communication from the severe interference generated by the TN. At the same time, the spatial spectrum sensing done by the NTN user helps manage the interference produced by the NTN towards the TN.

### v2

@@ Abstract
With the construction of low-earth orbit (LEO) satellite constellations, ubiquitous connectivity has been achieved. However, to accommodate the increasing communication needs in non-terrestrial networks (NTNs), wireless users in remote areas may require access to more spectrum than is currently allocated. Terrestrial networks (TNs), such as cellular networks, are deployed in specific areas, but many underused licensed spectrum bands remain in remote areas. Therefore, bringing NTNs to a shared spectrum with TNs can improve network capacity under reasonable interference management. However, in satellite-terrestrial integrated networks (STINs), the comprehensive coverage of a satellite and the unbalanced communication resources of STINs make it challenging to manage mutual interference between NTN and TN effectively. This article presents the fundamentals and prospects of spectrum sharing (SS) in STINs by introducing four SS frameworks, their potential application scenarios, and technical challenges. Furthermore, advanced SS approaches related to interference management in STINs and performance metrics of SS in STINs are introduced. Moreover, a preliminary performance evaluation showcases the potential for sharing the spectrum between NTN and TN. Finally, future research opportunities for SS in STINs are discussed.
@@ I Introduction
Due to the boosting need for broader and more reliable wireless network coverage and data transmission, non-terrestrial networks (NTNs) have emerged as a promising solution. In urban areas with heavy traffic or remote areas where terrestrial networks (TNs) face geographical deployment limitations, NTNs can provide seamless access and ubiquitous connectivity [1]. Spectrum integration in satellite-terrestrial integrated networks (STINs) is a necessary step to improve the utilization efficiency of scarce spectrum resources. However, the coexistence of NTNs and TNs may be challenging, especially with regard to spectrum utilization and interference management. When NTN nodes such as low-earth orbit (LEO) satellites are incorporated into existing TNs via certain frequency bands, interference patterns can become more complicated due to different spatial propagation characteristics of signals. For example, the use of the Ka band or the Ku band for satellites may cause severe co-channel interference in existing communication systems. Moreover, within a fixed spectrum allocation, it is possible that while certain portions of the spectrum in NTNs are fully occupied, other segments within TNs may remain unutilized. The imbalance will lead to under-utilization of available spectrum resources and a significant hindrance to substantial improvements in system performance.
@@
In recent years, the rapid growth of NTNs has paved the way toward realizing STINs. STINs, a promising candidate for future sixth-generation (6G) wireless communications, incorporate LEO nodes into existing TN to take advantage of NTN and meet diverse communication requirements [2]. By using STINs, the coverage, flexibility, and spectrum efficiency of the wireless network can be significantly promoted, as terrestrial users can access various TN and NTN nodes from vast geographical areas. Inevitably, the spectrum resource will become scarcer in STIN as more NTN nodes are introduced. This encourages us to manage the spectrum when TN and NTN co-exist in the STIN [3]. Meanwhile, the licensed spectrum to the static TN nodes cannot always be utilized efficiently, as spatial spectrum holes cannot be eliminated within the geographic area of interest. Hence, mobile NTN users can admit access to users within the spectrum holes and promote spectrum efficiency [4]. In existing work, one of the promising technologies for TN and NTN coexistence is spectrum sharing (SS), which allows the NTN to share the spectrum pre-allocated to the TN. SS allows the NTN to work in the same frequency band when the TN users are active, as long as its interference to the TN can be tolerated [4, 5]. In this way, the STIN can provide high-quality and seamless wireless services at a slight cost of performance loss under the practical restrictions of limited spectrum resources.
@@
With the distinct advantages above, STINs can overwhelm conventional TNs in a variety of practical fields. As the key technique in STINs, SS becomes the main bottleneck in network performance promotion and faces several challenges [6]. First, a universal SS architecture cannot always meet the diverse needs of the STIN. Therefore, the SS frameworks for all possible STIN transmission scenarios should be carefully designed, including the uplink and downlink of both TN and NTN. Second, since interference in STIN leads to system performance degradation, mitigating interference becomes the most critical issue. Generally, such mitigation can be implemented by isolating users from the interference source and carefully managing the TN and NTN spectrum [7]. Meanwhile, the TN and NTN spectrum should be allocated appropriately, and the spectrum access scheme should be carefully designed [8]. In addition, since STINs are hierarchical and heterogeneous, the proper selection of performance metrics for system evaluation is also essential to the system, depending on the design objectives.
@@
To address these concerns, we first consider the isolation of STIN users. One of the well-known approaches for this is the utilization of a protection zone for TN, where NTN users within the protection zone are not allowed to access the spectrum shared by TN and NTN. However, the appropriate definition of protection zones still remains a challenging problem. Meanwhile, some studies have been carried out for spectrum allocation and management. Typical work involves spatial SS, dynamic SS, and game-theoretic SS approaches, which are potentially beneficial for STINs. In spatial SS, the objective is to find the geographic areas that are good for NTN access in STINs [9]. In dynamic SS, the objective usually becomes finding the temporal holes in the spectrum resource by treating the NTN as a secondary network [10]. Both spatial SS and dynamic SS can be considered as frequency reuse schemes in STINs. For spectrum allocation and sharing, game models are competitive candidates in finding the optimal scheme by modeling TN and NTN as game players [11]. All of the above topics can be further studied to meet the performance requirements of future STINs.
@@
This article focuses on the SS frameworks for various STIN scenarios and the state-of-the-art approaches to SS in STINs. The main contributions are summarized as follows.
@@
  • We summarize the potential performance metrics for evaluating the efficiency of SS in STINs and validate the improvement via a practical case study. Finally, future research opportunities for SS in STINs are highlighted.
@@ II Frameworks of SS in STINs
In this section, we present four frameworks of SS in STINs as illustrated in Fig. 2. By implementing SS for NTN and TN in STINs, the total bandwidth is divided into two segments: 1) shared spectrum for both TN and NTN, and 2) reserved spectrum exclusively for NTN, as shown at the top of Fig. 2.
@@ II-A NTN Downlink and TN Downlink SS
When a satellite transmits signals to an NTN user and a BS transmits signals to a TN user, they can access the shared spectrum, as shown in Fig. 2(1). For example, this framework corresponds to the NTN satellite n256 frequency band and new radio (NR) n65 and n66 frequency bands in the International Telecommunication Union Radiocommunication Sector (ITU-R) and the 3rd Generation Partnership Project (3GPP) [12]. It’s worth noting that SS methods not only apply to the current spectrum regulations but also facilitate future spectrum utilization and standards development. In this case, TN users can be regarded as primary users and NTN users as secondary users. Designing a three-dimensional (3D) protection zone for each TN user is essential. The protection zone can be either spherical or tapered. Satellites outside this protection zone may utilize the shared spectrum while those within continue to use the reserved spectrum. This is because sharing spectrum between TN and NTN will not only enhance the communication performance of TN users through the expanded bandwidth but also result in interference. The protection zone is mainly set to protect the quality of service (QoS) of TN users. Theoretically, NTN users can also be affected by interference leakage from BSs, considering their beam directionality on the antenna surface in the horizontal plane. However, near-vertical transmissions from satellites can hardly be influenced. Moreover, satellites can mitigate mutual interference by constraining spectrum access within a specified distance range.
@@ II-B NTN Downlink and TN Uplink SS
Consider a scenario involving sharing a reverse link spectrum in which an NTN user receives signals from a satellite. As depicted in Fig. 2(2), both the satellite and the NTN user can access the shared spectrum through TN uplink communications. In this framework, satellite spectrum sensing can be performed to detect the strength of TN uplink signals to reduce mutual interference. If the uplink signal exhibits a low level of strength, the NTN downlink transmission can be operated to prevent significant mutual interference. To be more specific, due to the directionality of satellite downlink signals, the corresponding interference to TN uplink transmissions on the near-horizontal plane is relatively small. In addition, owing to the extended propagation distance and limited antenna gain associated with the satellite’s side lobes, TN uplink users located within the reception range of these side lobes exert minimal interference on the satellites’ spectrum sensing capabilities. In particular, spectrum sensing is conducted because satellites and NTN users can neither determine TN users transmitting in the current time slot nor predict aggregated interference.
@@ II-C NTN Uplink and TN Downlink SS
When an NTN user transmits signals to a satellite, they can opportunistically access the shared spectrum with TN downlink communications, as shown in Fig. 2(3). For example, this framework corresponds to the NTN satellite n256 frequency band and NR n2, n25, and n70 frequency bands in ITU-R. This framework is feasible due to a couple of factors. First, the BS is tilted downward for TN downlink communications and employs beamforming techniques to serve TN users. Because of the BS’s antenna angle and beamforming, interference leakage from the BS to the satellite is minimal. Second, during NTN uplink communications, the NTN user and the satellite can collaboratively perform spectrum sensing. This allows them to opportunistically access the spectrum while minimizing interference to and from TN users. It is important to note that, based on spectrum sensing conducted by NTN users, the aggregated interference power generated by NTN users is relatively lower than the desired signal power of the TN downlink communication. This is due to the high transmit power associated with TN BS, which also employs beamforming techniques.
@@ II-D NTN Uplink and TN Uplink SS
In the ITU-R, NTN satellite n256 frequency band and NR n65 frequency band overlap. As shown in Fig. 2(4), allowing NTN uplink communications to share the spectrum with TN uplink communications is facing many problems, which makes it necessary for NTN to use the n256 band in areas where TN does not use the n65 band. First, numerous TN uplink users operate within a satellite’s broad coverage area, and the distances between these users and NTN uplink users can be significant. This results in limited detection capabilities, making it difficult for either type of user to perceive the other effectively. Second, TN users typically have a limited number of antennas, leading to considerable interference leakage from TN users in the vertical direction. This interference can overwhelm the desired signals from NTN uplink communications. Third, NTN users are likely to experience a high rate of missed detections regarding the presence of TN users. Consequently, this inability to accurately sense TN users can lead to significant interference from NTN uplink transmissions to TN communications. As a result, TN users’ communication performance often falls short of expectations. Sharing the spectrum between these two co-directional uplink links is unfavorable from the perspective of interference and sensing capabilities in many situations. However, this spectrum-sharing framework may be viable with appropriate spectrum-aware strategies when the number of TN uplink users is relatively small.
@@ III-A Protection Zone-Based Approach
Method: Protection zones act as an auxiliary approach in SS as they define some geographical area in which users cannot access the spectrum of the primary system. SS techniques enable STIN users to sense and access the spectrum occupied by primary users to improve the spectrum efficiency. However, primary system interference cannot be ignored when users are close to interference sources. As protection zones are usually located around primary users, they can isolate interference for the users inside them, as shown in Fig. 3(a). While the users are outside the protection zones, they are free to access the spectrum.
Applications: With protection zones and knowledge of the user location in STINs, the SS technique enables intelligent monitoring of spatial spectrum holes and shares the spectrum outside the protection zone.
Difficulties and Solutions: Since the NTN is mobile and provides flexible spectrum access for users, finding the boundary for the protection zone becomes a challenging issue. The dynamics of STIN architectures also have significant impacts on the protection zones. Both challenges require statistical modeling and analysis based on user distributions.
III-B Spatial Spectrum Sensing (SSS)-Based Approach
@@
Method: SSS helps secondary users access the spectrum that relies on sensing rather than knowing the positions of interfering nodes [13]. The term “spatial” refers to the ability of this method to capture the spatial characteristics of large-scale networks. The process begins with evaluating the interference power for a secondary user seeking to access a spectrum occupied by primary users. As shown in Fig. 3(b), if the test statistic for the interference power exceeds a predefined energy detection threshold, it indicates that the channel is currently busy. In this case, the secondary user transmits with a low probability. Conversely, if the interference power is below the threshold, the secondary user transmits with a high probability.
Applications: When SSS is implemented in STINs, NTN-related devices act as secondary users to monitor the busy status of the shared channel. The SSS-based approach suits the NTN downlink and TN uplink SS framework. Acquiring the locations of currently transmitting TN users at NTN satellites is difficult in practical engineering. By selecting an appropriate energy detection threshold, we can manage the interference generated by NTN toward TN.
Difficulties and Solutions: The primary technical challenge of this method is accurately determining the probabilities of false alarms and missed detections. It is crucial to control these probabilities following the Neyman-Pearson criterion. Possible solutions include calculating probability distributions based on the statistical properties of the interference under various events or using machine learning techniques to fit these probabilities.
III-C Joint Spatial Spectrum Sensing (JSSS)-Based Approach
@@
Method: JSSS involves two different types of network nodes working together to perform SSS. As shown in Fig. 3(c), if both nodes determine that the channel is available for transmission, the secondary user will likely access the channel with a high probability. Conversely, if they disagree on availability, the secondary user will have a lower probability of accessing the channel.
Applications: This approach is particularly suitable for the NTN uplink and TN downlink SS framework. The reason is that the satellite, acting as the NTN receiver, may sometimes experience significant interference from the TN when the number of ground base stations is high in specific areas. Thus, the vertical interference leakage could not be ignored. As a result, the SSS performed by the satellite protects its communication from the interference generated by the TN. At the same time, the SSS done by the NTN user helps manage the interference produced by the NTN towards the TN.
Difficulties and Solutions: Since there are two different types of network nodes in JSSS, collaborating to decide whether SS can be accomplished has become a significant challenge. The energy detection thresholds of the satellite and the NTN user can be jointly optimized under the Neyman-Pearson criterion.